# Relaxed Equivariance via Multitask Learning

### v2

@@ 6.4 Computational Complexity
In this section, we report the computational time for the Geometric Algebra Transformer (GATr) and Transformer architectures. We selected models with an equivalent number of blocks and parameters for a fair comparison. GATr incorporates a unique design that includes a multivector parameter, while we adjusted the Transformer to match the parameter count of GATr. Both models have around $2.6$M parameters. Detailed configurations are provided in Appendix A.4. SE(3)-Transformer gives out of memory for this setting. We measured the computational efficiency of each model by recording the time taken for both forward and backward passes during training, as well as inference time. For the Transformer’s computations, we also considered all the cases of data augmentation and our training procedure with the equivariance loss (constant and gradual penalties). Figure 5 includes the wall-clock time as a function of batch size with a fixed number of nodes.
In all comparisons, GATr architecture consistently required the highest time, being approximately ten times slower than Transformer architecture. This significant difference can be attributed to the calculations of multivectors in GATr’s design. In the combined forward and backward passes, the addition of equivariance loss increases the computation time compared to the standard Transformer, due to the need to calculate two model outputs at each step. However, it’s still around $2.5$ times faster than GATr, in the worst case of a gradual penalty. Furthermore, GATr reached its memory capacity earlier, hitting an out-of-memory issue at a batch size of $2^{11}$. During inference, the computational speed for the Transformer trained with equivariance loss or data augmentation matches the standard Transformer, as all the differences applied in training. This results in an inference speed that is $10$ times faster than GATr.
6.5 Loss Surface
In this section, we analyze the relative ease of training equivariant models compared to non-equivariant models by examining the loss surface around the achieved local minima for each model. We explore how each architecture influences the loss landscape when trained on the same task. However, due to the high dimensionality of parameter spaces in neural networks, visualizing their loss functions in three dimensions might be a significant challenge. We use the filter normalization method introduced by Li et al. (2018), which calculates the loss function along two randomly selected Gaussian directions in the parameters space, starting from the optimal parameters $\theta^{*}$ achieved at the end of training.
@@ 7 Conclusion
We introduced a novel, simple method for learning approximate equivariance in a non-constrained setting through optimization. We formulated equivariance as a new weighted loss that is simultaneously optimized with the objective loss during the training process. We demonstrated that we can control the level of approximate equivariance based on the specific requirements of the task. Our method competes with or outperforms constrained equivariant baselines, achieving up to $10$ times faster inference speed and $2.5$ times faster training speed.
Limitations and Future Directions. While we showed that unconstrained models exhibit a more convex loss landscape near the local minima compared to equivariant models, this observation is subject to certain limitations. Specifically, we did not account for the trajectories that these models traverse to reach their respective minima. Understanding the optimization paths and how different initialization settings influence these paths remains unexplored. In future work, we aim to analyze the optimization process of each model and how it behaves during training.
@@ Acknowledgments
Appendix A Implementation Details
A.1 N-Body Dynamical System
@@
Following the methodology outlined in Brehmer et al. (2023), the dataset for the N-body system simulation encompasses four objects per sample. The center object is assigned a mass ranging from $1$ to $10$, whereas the other objects are uniformly positioned at a radius from $0.1$ to $1.0$ with masses between $0.01$ and $0.1$. We structured the datasets into two setups: in-distribution and out-of-distribution (OOD). Each sample in the in-distribution dataset is subjected to a random rotation within the range $[-10^{\circ},10^{\circ}]$. REMUL and data augmentation are trained with random rotations in the same range. Conversely, the OOD dataset is designed to evaluate the model’s generalization capabilities by incorporating extreme rotational perturbations, specifically with angles set within the ranges $[-180^{\circ},-90^{\circ}]$ and $[90^{\circ},180^{\circ}]$. We trained on $100$ samples, and each of the validation, test, and OOD datasets contains $5000$ samples. For models hyperparameters and training, we follow the same settings in Brehmer et al. (2023), summarized in Table 3. For REMUL, initial $\alpha=1$.
@@
#attention blocks $10$ $10$
#channels $128$ $8$ $384$
#attention heads $8$ $1$ $8$
#multivector $16$
#layers - $4$
#degrees - $4$
@@
A.2 Motion Capture
Motion Capture dataset by CMU (2003) features 3D trajectory data that records a range of human motions, and the task involves predicting the final trajectory based on initial positions and velocities. We have reported results for two types of motion: Walking (Subject #$35$) and Running (Subject #$9$).
@@ A.2 Motion Capture
#attention blocks $12$ $10$
#channels $128$ $8$ $384$
#attention heads $8$ $1$ $8$
#multivector $16$
#layers - $4$
#degrees - $4$
@@
#layers $3$ $3$ $3$ $3$
Following the standard experimental setup in the literature on this task (Han et al., 2022; Huang et al., 2022; Xu et al., 2024), we apply a train/validation/test split of $200$/$600$/$600$ for Walking and $200$/$240$/$240$ for Running. The interval between trajectories, $\Delta T=30$ for both tasks. For model hyperparameters, we fine-tuned around the same in Table 3 and found it the best for each model except for the Geometric Algebra Transformer we increased the attention blocks to $12$. We train each model for $2000$ epochs with batch size $=12$. For MLP comparisons, all models and baselines have the same number of layers and parameters. More details in Table 4. For REMUL, $\alpha=1$.
A.3 Molecular Dynamics
MD17 dataset (Chmiela et al., 2017) is a molecular dynamics benchmark that contains the trajectories of eight small molecules (Aspirin, Benzene, Ethanol, Malonaldehyde Naphthalene, Salicylic, Toluene, Uraci). We use the same dataset split in Huang et al. (2022); Xu et al. (2024), allocating $500$ samples for train, $2000$ for validation, and $2000$ for test. The interval between trajectories, $\Delta T=5000$. We selected the Equivariant Graph Neural Networks (EGNN) architecture and its non-equivariant version GNN, as introduced by Satorras et al. (2021). The input for GNN architecture is the initial positions along with atom types. Both architectures have the same hyperparameters, details in Table 5. For REMUL, $\alpha=1$.
@@ A.3 Molecular Dynamics
A.4 Computational Complexity
In the computational experiment of Geometric Algebra Transformer (GATr) and Transformer, we selected models with an equivalent number of blocks and parameters. GATr incorporates a unique design that includes a multivector parameter; we adjusted the Transformer architecture to match the parameter count of GATr. Both models have around $2.6$M parameters, detailed configurations are provided in Table 6. We selected a uniformly random Gaussian input with $20$ nodes and $7$ features dimension. We measured the computational efficiency of each model by recording the time taken for both forward and backward passes during training, as well as the inference time as a function of batch size. For each value, we took the average over $10$ runs with Nvidia A10101010 GPU.
@@ A.4 Computational Complexity
#attention heads $8$ $8$
#multivector $16$
Appendix B Additional Experiments
In this section, we present additional results on the three tasks: N-body dynamical system, Motion Capture, and Molecular Dynamics using the equivariance measure defined in Equation 13, which is consistent with our findings in the paper. We also include ablation studies on the number of samples required from the symmetry group during training for REMUL and data augmentation. Finally, we showcase molecules from the MD17 dataset, along with 2D and 3D visualizations of their structures.
B.1 N-Body Dynamical System
@@
Figure 6: N-body dynamical system. The second equivariance measure (defined in Equation 13). Plots from left to right: The first shows the Transformer trained with REMUL (gradual penalty), the second with a constant penalty, and the third presents the baselines (equivariant models, standard Transformer, and data augmentation). SE(3)-Transformer and GATr have a small equivariance error below the range of the plots ($3.1e^{-9}$ and $1.22e^{-14}$ respectively).
@@
Number of Samples from the Symmetry Group. We conduct ablation studies on the number of samples required from the symmetry group during training. We compare our training procedure, REMUL, with data augmentation method. We selected the N-body dynamical system with the same training details and hyperparameters indicated in Appendix A.1. As shown in Figure 7, REMUL achieves better performance using fewer samples from the symmetry group compared to data augmentation.
@@
Figure 7: Performance comparison of REMUL and data augmentation, using different numbers of samples from the symmetry group. Left: In-distribution performance and right: Out-of-distribution performance (N-body dynamical system).
B.2 Motion Capture

### v3

@@ 6.4 Computational Complexity
In this section, we report the computational time for the Geometric Algebra Transformer (GATr) and Transformer architectures. We selected models with an equivalent number of blocks and parameters for a fair comparison. Detailed configurations are provided in Appendix C.5. We measured the computational efficiency of each model by recording the time taken for both forward and backward passes during training, as well as inference time. For the Transformer’s computations, we also considered all the cases of data augmentation and our training procedure with the equivariance loss. Figure 3 includes the wall-clock time as a function of batch size with a fixed number of nodes.
(a) Combined forward pass
(b) Backward pass
(c) Inference time
Figure 3: Computational time for GATr and Transformer architectures. GATr has the highest time in all scenarios. Inference times for all versions of the Transformer (standard and trained with equivariance loss and data augmentation) are the same.
In all comparisons, GATr architecture consistently required the highest time, being approximately ten times slower than Transformer architecture. This significant difference can be attributed to the calculations of multivectors in GATr’s design. In the combined forward and backward passes, the addition of the equivariance loss increases the computation time of the standard Transformer as we calculate two model outputs at each step. However, it’s still around $2.5\times$ faster than GATr, in the worst case of a gradual penalty. Furthermore, GATr reached its memory capacity earlier, hitting an out-of-memory issue at a batch size of $2^{11}$. During inference, the computational speed for the Transformer trained with equivariance loss or data augmentation matches the standard Transformer, which results in an inference speed that is $10\times$ faster than GATr. Notably, while we include GATr as our equivariant baseline, GATr itself is computationally more efficient than many equivariant architectures such as SE(3)-Transformer and SEGNN, as indicated in [45].
@@ 7 Conclusion
[…]
Limitations and Future Directions. This work introduces a simple approach for understanding and analyzing unconstrained versus equivariant models, which significantly impact the field by enabling broader applicability and easier integration into existing frameworks. Building on these foundations, numerous additional ideas for extending our study present exciting opportunities for future research. For instance, as we noted earlier, $\alpha$ and $\beta$ serve as additional hyperparameters that could be constant or automatically updated with GradNorm algorithm, we could explore more efficient learnable weights, such as [97, 98]. Another promising avenue is applying our method during the fine-tuning phase when leveraging pre-trained models for tasks that require equivariance [99, 100]. On the other side, further analysis is required to understand the theoretical guarantees of approximate equivariance offered by REMUL, such as how relaxing equivariance constraints affects the model’s generalization bounds [47].
@@ Acknowledgments
Appendix A Loss Surface
In this section, we analyze the relative ease of training equivariant models compared to non-equivariant models by examining the loss surface around the achieved local minima for each model. We explore how each architecture influences the loss landscape when trained on the same task. However, due to the high dimensionality of parameter spaces in neural networks, visualizing their loss functions in three dimensions might be a significant challenge. We use the filter normalization method introduced by [101], which calculates the loss function along two randomly selected Gaussian directions in the parameters space, starting from the optimal parameters $\theta^{*}$ achieved at the end of training.
(a) Geometric Algebra Transformer
(b) Transformer
Appendix B Proofs
B.1 Propositions
Proposition 1. Let $f_{\alpha,\beta}\in\arg\min_{f\in\mathcal{H}}\widehat{\mathcal{L}}_{\mathrm{total}}(f;\alpha,\beta)$ be an empirical minimizer of the REMUL objective, and let $f^{\star}_{\mathrm{obj}}\in\arg\min_{f\in\mathcal{H}}\widehat{\mathcal{L}}_{\mathrm{obj}}(f)$ be an empirical minimizer for the objective loss alone. Then:
  1. (a)
    $f_{\alpha,\beta}$ is Pareto optimal for the bi-objective problem $(\min\widehat{\mathcal{L}}_{\mathrm{obj}}(f),\min\widehat{\mathcal{L}}_{\mathrm{equi}}(f))$.
  2. (b)
    The following trade-off inequality holds:
    $\widehat{\mathcal{L}}_{\mathrm{obj}}(f_{\alpha,\beta})-\widehat{\mathcal{L}}_{\mathrm{obj}}(f^{\star}_{\mathrm{obj}})\leq\frac{\beta}{\alpha}\left(\widehat{\mathcal{L}}_{\mathrm{equi}}(f^{\star}_{\mathrm{obj}})-\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\alpha,\beta})\right).$ (10)
Proof.
Let $f_{\alpha,\beta}$ be an empirical minimizer of $\widehat{\mathcal{L}}_{\mathrm{total}}(f;\alpha,\beta)$. By definition, for any $\tilde{f}\in\mathcal{H}$:
[…]
To show Eq. 10: Rearrange the optimality condition:
$\alpha(\widehat{\mathcal{L}}_{\mathrm{obj}}(f_{\alpha,\beta})-\widehat{\mathcal{L}}_{\mathrm{obj}}(\tilde{f}))\leq\beta(\widehat{\mathcal{L}}_{\mathrm{equi}}(f^{\prime})-\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\alpha,\beta})).$
Dividing by $\alpha>0$:
$\widehat{\mathcal{L}}_{\mathrm{obj}}(f_{\alpha,\beta})-\widehat{\mathcal{L}}_{\mathrm{obj}}(\tilde{f})\leq\frac{\beta}{\alpha}(\widehat{\mathcal{L}}_{\mathrm{equi}}(\tilde{f})-\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\alpha,\beta})).$
Setting $\tilde{f}=f^{\star}_{\mathrm{obj}}$ yields Eq. 10.
For Pareto Optimality: Assume, for contradiction, that $f_{\alpha,\beta}$ is not Pareto optimal. Then there exists an $\widetilde{\widetilde{f}}\in\mathcal{H}$ such that $\widehat{\mathcal{L}}_{\mathrm{obj}}(\widetilde{\widetilde{f}})\leq\widehat{\mathcal{L}}_{\mathrm{obj}}(f_{\alpha,\beta})$ and $\widehat{\mathcal{L}}_{\mathrm{equi}}(\widetilde{\widetilde{f}})\leq\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\alpha,\beta})$, with at least one of these inequalities being strict. Since $\alpha>0$ and $\beta>0$, this would imply $\alpha\widehat{\mathcal{L}}_{\mathrm{obj}}(\widetilde{\widetilde{f}})+\beta\widehat{\mathcal{L}}_{\mathrm{equi}}(\widetilde{\widetilde{f}})<\alpha\widehat{\mathcal{L}}_{\mathrm{obj}}(f_{\alpha,\beta})+\beta\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\alpha,\beta})$. This contradicts the assumption that $f_{\alpha,\beta}$ is a minimizer of $\widehat{\mathcal{L}}_{\mathrm{total}}(f;\alpha,\beta)$. Therefore, $f_{\alpha,\beta}$ must be Pareto optimal. ∎
B.2 Equivariance Measure
We define the equivariance metric $E$ to quantify the degree of equivariance exhibited by a function $f$, as:
$E(f,G)=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\left\|\frac{1}{M}\sum_{i=1}^{M}\rho(g_{i})(f(x))-\frac{1}{M}\sum_{i=1}^{M}f(\phi(g_{i})(x))\right\|_{2}$ (11)
Proof.
Starting from Eq. 1: $f(\phi(g)(x))=\rho(g)(f(x)$, the group integration of both sides w.r.t. the normalized Haar measure $\mu$ yields:
$\int_{G}f(\phi(g)(x))\,d\mu(g)=\int_{G}\rho(g)(f(x))\,d\mu(g)$ (12)
$\int_{G}f(\phi(g)(x))\,d\mu(g)\approx\frac{1}{M}\sum_{i=1}^{M}f(\phi(g_{i})(x))$ (13)
$\int_{G}\rho(g)(f(x))\,d\mu(g)\approx\frac{1}{M}\sum_{i=1}^{M}\rho(g_{i})(f(x))$ (14)
where $M$ is a large number of samples from $G$.
Given the group averages, we can then define the equivariance error $E(f,G)$ as the average norm of the difference between these two averages over the data distribution $\mathcal{D}$:
$E(f,G)=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\left\|\frac{1}{M}\sum_{i=1}^{M}\rho(g_{i})(f(x))-\frac{1}{M}\sum_{i=1}^{M}f(\phi(g_{i})(x))\right\|_{2}$ (15)
with $\|\cdot\|_{2}$ denotes an $L_{2}$ norm (for non-scalar function).
Appendix C Implementation Details
C.1 Equivariance Loss
The empirical equivariance loss defined in Eq. 6, $\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\theta})=\tfrac{1}{n}\sum\mathbb{E}_{g\sim G}[\ell(f_{\theta}(\phi(g)x_{i}),\rho(g)y_{i})]$, measures the consistency of the model’s predictions on transformed inputs against the correspondingly transformed ground truth labels. It is distinct from a direct measure of functional equivariance, which compare $f_{\theta}(\phi(g)x_{i})$ with $\rho(g)f_{\theta}(x_{i})$ (the transformed prediction of the original input). While the latter directly assesses the equivariance of the function $f_{\theta}$ itself, our choice of $\widehat{\mathcal{L}}_{\mathrm{equi}}$ offers a crucial advantage: it continuously anchors the learning process to the ground truth. To see this, let $f_{\theta}(x)=y(x)+\gamma(x)$, where $y(x)$ is the true label for input $x$ and $\gamma(x)$ is the model’s prediction error. If we assume the ground truth data itself is perfectly equivariant, i.e., $y(\phi(g)x)=\rho(g)y(x)$, then the term minimized by $\widehat{\mathcal{L}}_{\mathrm{equi}}$ (for a single instance, taking $\ell$ as an $L_{p}$ norm) becomes:
$\|\underbrace{f_{\theta}(\phi(g)x_{i})}_{y(\phi(g)x_{i})+\gamma(\phi(g)x_{i})}-\underbrace{\rho(g)y_{i}}_{y(\phi(g)x_{i})}\|_{p}=\|\gamma(\phi(g)x_{i})\|_{p}.$
Thus, minimizing $\widehat{\mathcal{L}}_{\mathrm{equi}}$ directly minimizes the magnitude of the prediction error on transformed inputs. This helps prevent the model from "drifting" into solutions that might be equivariant but incorrect (i.e., $f_{\theta}(\phi(g)x_{i})\approx\rho(g)f_{\theta}(x_{i})$ but both are far from $\rho(g)y_{i}$). In contrast, a loss term based on functional equivariance, $\|f_{\theta}(\phi(g)x_{i})-\rho(g)f_{\theta}(x_{i})\|_{p}$, would simplify to $\|\gamma(\phi(g)x_{i})-\rho(g)\gamma(x_{i})\|_{p}$. While this term directly encourages the *error itself* to be equivariant, minimizing it alone does not guarantee that the error magnitude $\|\gamma(\cdot)\|_{p}$ is small. Our REMUL objective, by combining $\alpha\widehat{\mathcal{L}}_{\mathrm{obj}}(f_{\theta})$ (which minimizes $\|\gamma(x_{i})\|_{p}$ on original data) with $\beta\widehat{\mathcal{L}}_{\mathrm{equi}}(f_{\theta})$ (which minimizes $\|\gamma(\phi(g)x_{i})\|_{p}$ on transformed data, given ideal data equivariance), aims for both accuracy and consistency under transformations. The degree to which this also induces functional equivariance in $f_{\theta}$ (i.e., making $\|\gamma(\phi(g)x_{i})-\rho(g)\gamma(x_{i})\|_{p}$ small) is then assessed empirically using the equivariance metrics $E$ and $E^{\prime}$ as shown in our experiments.
C.2 N-Body Dynamical System
@@
Following the methodology outlined in [45], the dataset for the N-body system simulation encompasses four objects per sample. The center object is assigned a mass ranging from $1$ to $10$, whereas the other objects are uniformly positioned at a radius from $0.1$ to $1.0$ with masses between $0.01$ and $0.1$. We structured the datasets into two setups: in-distribution and out-of-distribution (OOD). Each sample in the in-distribution dataset is subjected to a random rotation within the range $10^{\circ},10^{\circ}]$. REMUL and data augmentation are trained with random rotations in the same range. Conversely, the OOD dataset is designed to evaluate the model’s generalization capabilities by incorporating extreme rotational perturbations, specifically with angles set within the ranges $180^{\circ},-90^{\circ}]$ and $90^{\circ},180^{\circ}]$. We trained on $100$ samples, and each of the validation, test, and OOD datasets contains $5000$ samples. For models hyperparameters and training, we follow the same settings in [45], summarized in Table 4. For REMUL, initial $\alpha=1$.
@@
#attention blocks $10$ - $10$
#channels $128$ 88 $384$
#attention heads 88 11 88
#multivector $16$ - -
#layers - 44 -
#degrees - 44 -
@@
C.3 Motion Capture
Motion Capture dataset by [91] features 3D trajectory data that records a range of human motions, and the task involves predicting the final trajectory based on initial positions and velocities. We have reported results for two types of motion: Walking (Subject #$35$) and Running (Subject #$9$).
@@ C.3 Motion Capture
#attention blocks $12$ - $10$
#channels $128$ 88 $384$
#attention heads 88 11 88
#multivector $16$ - -
#layers - 44 -
#degrees - 44 -
@@
#layers 33 33 33 33
Following the standard experimental setup in the literature on this task [3, 92, 4], we apply a train/validation/test split of 200200/600600/$600$ for Walking and 200200/240240/$240$ for Running. The interval between trajectories, $\Delta T=30$ for both tasks. For model hyperparameters, we fine-tuned around the same in Table 4 and found it the best for each model except for the Geometric Algebra Transformer we increased the attention blocks to $12$. We train each model for $2000$ epochs with batch size =12=12. For MLP comparisons, all models and baselines have the same number of layers and parameters. More details in Table 5. For REMUL, $\alpha=1$.
C.4 Molecular Dynamics
MD17 dataset [95] is a molecular dynamics benchmark that contains the trajectories of eight small molecules (Aspirin, Benzene, Ethanol, Malonaldehyde Naphthalene, Salicylic, Toluene, Uraci). We use the same dataset split in [92, 4], allocating $500$ samples for train, $2000$ for validation, and $2000$ for test. The interval between trajectories, $\Delta T=5000$. We selected the Equivariant Graph Neural Networks (EGNN) architecture and its non-equivariant version GNN, as introduced by [5]. The input for GNN architecture is the initial positions along with atom types. Both architectures have the same hyperparameters, details in Table 6. For REMUL, $\alpha=1$.
@@ C.4 Molecular Dynamics
C.5 Computational Complexity
In the computational experiment of Geometric Algebra Transformer (GATr) and Transformer, we selected models with an equivalent number of blocks and parameters. GATr incorporates a unique design that includes a multivector parameter; we adjusted the Transformer architecture to match the parameter count of GATr. Both models have around 2.62.6M parameters, detailed configurations are provided in Table 7. SE(3)-Transformer gives out of memory for this setting. We selected a uniformly random Gaussian input with $20$ nodes and $7$ features dimension. We measured the computational efficiency of each model by recording the time taken for both forward and backward passes during training, as well as the inference time as a function of batch size. For each value, we took the average over $10$ runs with Nvidia A1010 GPU.
@@ C.5 Computational Complexity
#attention heads 88 88
#multivector $16$ -
C.6 Compute Resources
We ran all the experiments using a single Nvidia A10 GPU.
Appendix D Additional Experiments
This section provides additional experimental results to further validate our training procedure REMUL. We include:
  • Additional evaluations on the three tasks: For the N-body dynamical system, Motion Capture, and Molecular Dynamics (MD17), we include the standard equivariance error $E^{\prime}$ (defined in Eq. 9 of the main paper). These results are consistent with our findings in the paper, using the $E$ metric.
  • Performance on MD17: We include detailed results showing performance and equivariance error trade-offs for REMUL applied to GNN architecture on individual molecules from the MD17 dataset (complementing Table 3 in the main paper).
  • Ablation on Group Sampling: We conduct an ablation study investigating the impact of the number of group samples used during training for REMUL and data augmentation.
  • Convergence Speed Analysis: We compare the convergence speed of REMUL against data augmentation by tracking training and validation MSE as a function of training steps.
  • Axis-Specific Equivariance Error on Motion Capture: To further investigate the nature of symmetries in the Motion Capture dataset, we report the equivariance error around individual $X$, $Y$, and $Z$ axes, separately.
  • Additional N-Body System Benchmark: We evaluate REMUL (applied to a GNN model) on an additional N-body system benchmark, comparing it against several equivariant architectures.
  • Molecular Structures: We provide 2D and 3D visualizations of the molecules from the MD17 dataset.
D.1 N-Body Dynamical System
(a) REMUL: Gradual penalty
$0.01$
$0.1$
$1.0$
$10.0$
$100.0$
(b) REMUL: Constant penalty
(c) Baselines
@@ D.1 N-Body Dynamical System
Figure 5: N-body dynamical system. The second equivariance measure $E^{\prime}$. (a) Transformer trained with REMUL (gradual penalty), (b) Transformer trained with REMUL (constant penalty), and (c) Baselines: Equivariant models, standard Transformer, and data augmentation.
D.2 Number of Samples from the Symmetry Group
@@
We conduct ablation studies on the number of samples required from the symmetry group during training. We compare our training procedure with data augmentation method. We selected the N-body dynamical system with the same training details and hyperparameters indicated in Appendix C.2. As shown in Figure 6, REMUL achieves better performance using fewer samples from the symmetry group compared to data augmentation.
(a) MSE: In-distribution
(b) MSE: Out-of-distribution
@@
Figure 6: Performance comparison of REMUL and data augmentation on N-body dynamical system, using different numbers of samples from the symmetry group.
D.3 Convergence Speed Analysis
To assess the training efficiency of REMUL relative to data augmentation, we analyzed their convergence behavior on the N-body dynamical system. Both REMUL and DA were applied to a standard Transformer using the same experimental settings described in Appendix C.2. We report the Mean Squared Error (MSE) of the training and validation samples as a function of the training steps. The results, presented in Figure 7, indicate that REMUL achieves lower training and validation errors compared to data augmentation.
(a) MSE: Training samples
(b) MSE: Validation samples
Figure 7: Comparison of convergence speed between REMUL and data augmentation on the N-body system.
D.4 Additional Benchmark
To further assess REMUL’s capabilities, we evaluated it on the N-body system benchmark from [5]. We applied REMUL to a Graph Neural Network (GNN) architecture, using the same hyperparameter configurations in [5] for a fair comparison. Table 8 compares our approach against several equivariant models, including EGNN [5], SEGNN [6], FA-GNN [24], and TFN [14]. The results demonstrate that REMUL achieves strong performance, outperforming EGNN and FA-GNN while being competitive with SEGNN, despite the latter incorporates more specialized geometric features.
Table 8: Additional benchmark on N Body system.
MSE
SE(3)-Tr 0.0244
TFN 0.0155
MPNN 0.0107
EGNN 0.0071
SEGNN 0.0043
FA-GNN 0.0057
REMUL-GNN 0.0046
D.5 Motion Capture
(a) Equiv. error: Walking
(b) Equiv. error: Running